OpenStudy (anonymous):

how do you get 2^(-7)=x^(5)

OpenStudy (anonymous):

2\[2^{-7}=x^{5}\]

OpenStudy (anonymous):

I don't think I'm understanding the problem, do you want to know how to solve for x?

OpenStudy (anonymous):

yea

OpenStudy (jdoe0001):

hmmm get teh fifth root for each side

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

You can introduce a new exponent as long as it is done to each side. For example, (x^5)^1/5=(2^-7)^1/5

OpenStudy (anonymous):

What would you get when you simplify that?

OpenStudy (anonymous):

I don't know lol

OpenStudy (jdoe0001):

\(\bf 2^{-7}=x^5\implies \cfrac{1}{2^7}=x^5\implies \sqrt[5]{\cfrac{1}{2^7}}=\sqrt[5]{x^5}\implies \cfrac{\sqrt[5]{1}}{\sqrt[5]{2^7}}=x \\ \quad \\ \cfrac{1}{\sqrt[5]{2^7}}=x\implies \square ?\)

OpenStudy (jdoe0001):

recall that \(\large a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\)

OpenStudy (anonymous):

That is basically what needs to be done, now do you know how the 5th root part simplifies at the end? What I'm trying to say is do you know how to turn that into one exponent?

OpenStudy (anonymous):

the answer at the end is 0.3789

OpenStudy (anonymous):

I didn't check but that's not what I was asking. Write out that 5th root of 2^7 as just 2 to some power

OpenStudy (anonymous):

This will be good practice later on when you want to keep exponents in place to make further simplification easier.

OpenStudy (anonymous):

I don't understand. If you can use the equation bar that will help.

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

\[\sqrt[5]{2^{7}} = ?\]

OpenStudy (anonymous):

Write this as 2 to some exponent

OpenStudy (anonymous):

4 but that's not the answer

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (anonymous):

I know it's not and I feel bad for dragging this out but it will help to know these basic exponent rules

OpenStudy (anonymous):

I know the basic exponent rules but what Jdoe did up their will not give me the correct answer

OpenStudy (anonymous):

1/5square(2^7) equals 1/4 which is not the answer

OpenStudy (anonymous):

I think jdoe is right though, \[1/\sqrt[5]{2^{7}}\]= x

OpenStudy (anonymous):

it is nvm

OpenStudy (anonymous):

Step by step: 2^-7=x^5 1/(2^7)=x^5 (1/(2^7))^1/5=x 1/(2^7)^1/5=x